OpenStudy (anonymous):

Chain rule help? Can anyone give examples of finding the derivative using the chain rule...

OpenStudy (anonymous):

find the derivative of \[\sqrt{1-x^2}\]

OpenStudy (anonymous):

or try this, first find the derivative of \[\frac{1-x}{1+x}\] and then use the chain rule to find the derivative of \[\frac{1-\sin(x)}{1+\sin(x)}\]

OpenStudy (anonymous):

Honestly, I have no idea how to do that.. Do you have to use the quotient rule first? Since this problem involves division?